OpenStudy (anonymous):

Solve the system: 22x-19y=28. 55x-29y=107.

OpenStudy (anonymous):

Can multiply first equation by 5 and second equation by -2 and add them together. That will eliminate the x variable and you can find y.

OpenStudy (anonymous):

Here's the setup: 5[22x-19y=28] ----> 110x-95y=140 -2[55x-29y=107] -> -110x+58y=-214 When you add the two equations together, x disappears and you can get a value for y. Plug that into the first equation (or the second, if you prefer) to find x.